# Entangelment Entropy on Generalized Brillouin Zone

### v1

@@ Abstract
We investigate the entanglement properties of non-Hermitian Su-Schrieffer-Heeger (SSH) model from the perspective of the Generalized Brillouin Zone (GBZ). The non-Bloch entanglement entropy is defined on a quasi-reciprocal lattice, obtained by performing an ordinary Fourier transformation on the non-Bloch Hamiltonian. We demonstrate that the broken bulk-boundary correspondence is recovered in terms of the non-Bloch entanglement entropy. When the GBZ is circular, we show that the non-Bloch entanglement entropy is well-defined (real and positive-definite) in large parameter regions, except close to the exceptional points (EPs). In the critical region, we found that each Fermi point contributes precisely 1 to the central charge $c$ of the logarithmic scaling. At the EP, the central charge becomes negative due to the presence of the exceptional bound state. For the case of non-circular GBZ, long-range hopping emerges in the quasi-reciprocal lattice, and the von Neumann entropy on the GBZ is no longer real. However, the non-Bloch edge entanglement entropy remains real, which serves as a reliable topological indicator and respects the bulk-boundary correspondence. We compute the topological phase diagram, and reveal the critical behavior along the exceptional phase boundaries.
@@
Non-Hermitian quantum mechanics naturally arises from the study of open quantum systems coupling with an external environment [1, 2]. For non-Hermitian systems with nonreciprocal hoppings, the bulk eigenstates under open boundary conditions (OBC) accumulate at boundaries, which is named by the non-Hermitian skin effect (NHSE) [3, 4, 5, 6] and has been extensively studied recently [7, 8]. In the presence of NHSE, the energy spectrum sensitively depends on the boundary conditions: the energy spectrum under OBC forms a star-shaped branching figures [9], while the energy spectrum under periodic boundary conditions (PBC) makes up loops enclosing the OBC spectrum [10, 11]. The NHSE also results in the breakdown of the celebrated bulk-boundary correspondence (BBC) of topological band theory [12, 13]: the transition points of bulk topological invariants do not match with transition points of zero modes under OBC [14, 15, 3]. In order to understand the topology of non-Hermitian systems, the non-Bloch band theory and the generalized Brillouin zone (GBZ) are proposed, based on which the BBC for non-Hermitian systems can be reestablished [3, 16, 17]. Motivated by this breakthrough, the non-Bloch band theory has been widely used, e.g., to construct Green function [18], to investigate the dynamical phenomena [19] and the non-Bloch parity-time symmetry [20].
@@
The entanglement properties of the non-Hermitian quantum systems is also an interesting topic which has attracted many attentions recently [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. The conception of EE has been generalized to the non-Hermitian systems in the biorthogonal basis [29, 28]. A logarithmic scaling with central charge $c=-2$ is found at critical point of the non-Hermitian Su-Schrieffer-Heeger (SSH) model, which can be explained by the $bc$-ghost non-unitary CFT [29, 38]. However, the physical meaning of EE remains somewhat ambiguous in this straightforward generalization [38, 40]: The eigenvalues of reduced density matrix $\rho_{A}$ are no longer positive definite, which leads to a negative or even complex EE. This is inconsistent with the probability interpretation of a measurable quantity [38, 40]. Some attempts have been made to remedy this situation, for example, the generic entanglement entropy $S_{A}=-\mathrm{Tr}(\rho_{A}\ln|\rho_{A}|)$ is proposed in Ref. [38], the modified partial trace formalism is introduced in Ref. [27]. However, more systematic analysis is needed in this direction.
Inspired by the success of non-Hermitian topological band theory in terms of GBZ, we may ask whether it is possible to construct an EE on the GBZ which has proper behaviors? We address this question in this paper for the non-Hermitian systems with skin effect. We begin with a GBZ Hamitonian and then make a Fourier transformation to obtain a quasi-reciprocal lattice [48]. The entanglement entropy on GBZ is calculated by partitioning this quasi-reciprocal lattice. The behaviors of various EE will be examined for two types of GBZ (circular or non-circular). Our study may shed light on the physical meanings of GBZ and non-Hermitian EE, and also open the possibility to study the many-body physics based on GBZ.
This paper is organized as the following: In Sec. II, we review the basic notions of biorthogonal entanglement entropy. In Sec. III, we introduce the non-Hermitian SSH model and the quasi-reciprocal lattice, on which the entanglement entropy is calculated. In Sec. IV, we present the results of entanglement entropy on the circular GBZ. In Sec. V, we present the results of entanglement entropy on the non-circular GBZ. Finally, a brief summary and discussion are provided in Sec. VI.
@@
in which the sites $i,j$ are restricted inside the subsystem $A$ and the index $m$ labels the occupied energy levels. The single-particle entanglement spectrum is obtained from the eigenvalue spectrum $\{\xi_{l}\}$ of the matrix $C^{A}$ under periodic boundary condition. Based on this, the von-Neumann and Rényi entropy are given by
@@
where $S^{\alpha}_{\text{OBC}}$ and $S^{\alpha}_{\text{PBC}}$ are $\alpha$-th Rényi entropies calculated under OBC and PBC, respectively. The edge entanglement entropy is a useful tool to extract the topological information of edge states [54, 30, 50].
To study the critical behavior of non-Hermitian systems, the entanglement entropy can be fitted using the following universal formula [55, 56],
@@ III Quasi-reciprocal Hamiltonian based on generalized Brillouin zone
in which $k\in[0,2\pi)$ being the crystal momentum. The hopping processes of this model in real space are illustrated in Fig. 1(a). According to the theory of GBZ [3, 16, 17], we substitute $e^{ik}$ with $\beta=e^{ik^{\prime}}$ ($k^{\prime}=k+i\tau(k)\in\mathbb{C}$) to gain the corresponding non-Bloch Hamiltonian
@@
Two typical GBZs are shown in Fig. 2(a) and (b), where the GBZ forms a closed loop encircling the origin on the complex plane. The OBC spectrum in the thermodynamic limit can be exactly obtained when $\beta$ goes along the loop of GBZ.
To define the entanglement entropy in terms of GBZ, one need to construct an artificial real-space lattice which respects the translation invariance. Suppose that the phase angle of $\beta_{GBZ}$ is equally distributed on the GBZ, we then construct non-Bloch Hamiltonian matrix $H(\beta_{GBZ})$ by selecting the points whose phase angle belongs to the Brillouin zone (BZ) of original physical lattices, that is,
@@
where $N$ is the number of unit cells of the lattice. The non-Bloch operator $c_{\beta(k_{m})}^{\dagger}$ can be transformed into a new lattice operator $\tilde{c}_{j}$ by a Fourier transformation,
@@
Figure 2: The examples of circular GBZ (a) and non-circular GBZ (b). The green dashed circle denotes the conventional BZ with $|\beta|=1$. The radius of circular GBZ (orange circle) in (a) is everywhere smaller than $1$, indicating that wave functions only accumulate at the left end of the chain. Sub-figures (c) and (d) show the energy spectra of original OBC Hamiltonian $H_{OBC}$ (blue dots) and quasi-reciprocal OBC Hamiltonian $\widetilde{H}_{OBC}$ (red circles). The length of lattice is set to be $L=40$. For the case of circular GBZ (c), the blue dots precisely match the red circles. However, for the non-circular GBZ (d), the blue dots slightly deviate from the red circles, especially close to the bifurcations. We fix $t_{3}=0$ in (a) and (c), but $t_{3}=0.2$ in (b) and (d). Other parameters used in the calculations are $t_{1}=0.4$, $t_{2}=1$, and $\gamma=2/3$.
@@
$\{\epsilon_{OBC}\}\stackrel{{\scriptstyle N\to\infty}}{{=}}\{\widetilde{\epsilon}_{OBC}\}\stackrel{{\scriptstyle N\to\infty}}{{=}}\{\widetilde{\epsilon}_{PBC}\}$ (19)
When next-nearest-neighbor (NNN) hoppings are included in the non-Hermitian SSH model, the degree of polynomial $\det[E-H(\beta)]$ is greater than 2, and the GBZ will become non-circular [3]. In this case, the module $|\beta|$ is not a constant but depends on the phase angle $k$. By Fourier transformation, this leads to the long-range hoppings in quasi-reciprocal lattices [48]. In practical calculations, we adopt an approximation to obtain the open boundary of quasi-reciprocal lattice: we neglect all the hopping terms whose hopping range is larger than half of the lattice length ($N/2$). This is reasonable since long-range hoppings are generically power-law decaying [48].
@@
with $\tilde{S}=\mathrm{diag}\{r,r,r^{2},r^{2},\cdots,r^{N-1},r^{N},r^{N}\}$. This transformation differs from the well-studied similarity transformation $\bar{S}=\mathrm{diag}\{1,r,r,r^{2},\cdots,r^{N-1},r^{N}\}$ [3], which renormalizes the intracell hopping amplitudes; see Fig. 1(c).
If $t_{3}\neq 0$, the long-range hoppings appears in the quasi-reciprocal lattices due to the non-constant $|\beta|$. There are tiny difference between the energy spectra of $H_{OBC}$ and $\widetilde{H}_{OBC}$ [see Fig. 2(d)], for the long-range hoppings can not be fully taken into account when the length of lattices is finite. The difference in energy spectra will decrease to zero in the thermodynamic limit.
@@ IV Entanglement entropy on Circular GBZ
We first present the results of circular GBZ in Fig. 3, in which we use the non-Hermitian SSH model with $t_{3}=0$ as an example. It is well-known that the BBC is broken in the non-Hermitian SSH model: the topological transition of zero-energy modes under OBC can not be explained by the Bloch Hamiltonian $h(k)$ under PBC [3, 16]; see Fig. 3(a). The entanglement entropy can be calculated by applying the correlation matrix method described in Sec. II to the original non-Hermitian Hamiltonian under PBC [28, 29]. We show the results of entanglement entropy for SSH model with $t_{3}=0$ in Fig. 3(b). The entanglement entropy is well-defined in the line-gap region, i.e., the values are real and positive-definite (away from the two exceptional points). However, in the point-gap region, the entanglement entropy is not well-defined: it becomes complex with a non-zero imaginary part [28, 50]. Noted that at the special point with $t_{1}=t_{2}$ in the point-gap region, the entanglement entropy is also real, as studied in Ref. [36]. In all, the entanglement entropy defined on the original PBC lattice basically reflects the characteristics of the Bloch band, and does not respect the BBC.
@@
The entanglement entropy drops to $2\log(2)$ at the point $t_{1}=\gamma$; see Fig. 3(c). At this point, the GBZ collapses into a single point with $r=0$, and the intra-(inter-)unit cell hoppings of the quasi-reciprocal lattice become $2\gamma$ and 0 (0 and $\infty$), respectively. By performing the similar transformation $S=\mathrm{diag}\{1,1/r,1,1/r,\cdots,1,1/r\}$, this case can be mapped into the dimerized limit with intra-unit cell hopping being zero; see Fig. 1(c). There are two isolated edge states, hence the entanglement entropy is exactly $2\log(2)$ [58]. Note, when $t_{1}<\gamma$, the Hamiltonian $H$ cannot be transformed into an Hermitian SSH model [3]. However, the entropy remains well-defined, as demonstrated in Fig. 3(c); see the Appendix A for the analysis of realness. This is a significant finding of the present paper.
@@
The quasi-reciprocal lattice Hamiltonian has an EP at $\gamma_{c}=\sqrt{t_{1}^{2}+t_{2}^{2}}$. We then turn to study the possible critical behavior in this model. In Fig. 4(a), the fitted $c$ is plotted as a function of $\gamma$ for $t_{1}=0.5$. For $\gamma<\gamma_{c}$, the value of $c$ is zero, indicating that the system is noncritical in this region. However, for $\gamma>\gamma_{c}$, the value of $c$ saturates to 2, implying critical behavior; See Fig. 4(b) for the representative logarithmic scaling of entanglement entropy with $L_{A}/L=1/2$. The critical behavior is associated with the Fermi surface structure of the energy spectrum [36]. Since we construct the ground state according to the real part of energy, the system is an insulator without any Fermi points when $\gamma<\gamma_{c}$, but when $\gamma>\gamma_{c}$ it becomes a gapless metal with two Fermi points. We therefore conclude that each Fermi point contributes exactly 1 to the central charge $c$ of the logarithmic scaling. This is different from the original SSH model (PBC) studied in Ref. [36], where each Fermi point contributes 1/2 to the value of $c$.
@@
The entanglement entropy diverges at the EP with $\gamma=\gamma_{c}$. The insets in Fig. 4(c) shows the two exceptional momentum $k_{EP}$ (0 and $\pi$) in this model. To avoid divergence, the entanglement entropy can be calculated by slightly shifting from the EP $\gamma=\gamma_{c}-\eta$ [29, 38]. Note that the dispersion of this EP is square root dependence on $k$ instead of linear. The scaling behavior of entanglement entropy for a fixed system length $L=400$ is shown in Fig. 4(d), where a tiny shift $\eta=1\times 10^{-8}$ is used and $L_{A}$ is chosen to include an odd number of unit cells. The extracted central charge is negative (around $-3$), which is attributed to the exceptional bound state close to EP [37]. However, the value of central charge slightly increases when the system length is increasing, as depicted in Fig. 4(c).
@@
Figure 4: (a) The central charges of entanglement entropy (real) on circular GBZ as a function of $\gamma$. The two insets display representative energy spectra on two sides of the EP $\gamma_{c}=\sqrt{1.25}$, where the solid (dashed) line indicate the occupied (unoccupied) states. (b) The logarithmic scaling of entanglement entropy (fixed ratio $L_{A}/L=1/2$) when $\gamma$ is far away from EP. (c) The fitted central charge using Eq. (10) versus total system length $L$. In the calculations, $\gamma$ is shifted from $\gamma_{c}$ by a tiny quantity $\eta=1\times 10^{-8}$. Insets display the energy spectrum and the dispersion relation, revealing two exceptional momenta at $k=0,\pi$ that render the momentum-space Hamiltonian defective. (d) The entanglement entropy as a function of the subsystem length $L_{A}$ for fixed total length $L=400$. The hopping parameters for all sub-figures are $t_{1}=0.5$, $t_{2}=1$, and $t_{3}=0$.

### v2

@@ Abstract
We investigate the entanglement properties of non-Hermitian Su-Schrieffer-Heeger (SSH) model from the perspective of the Generalized Brillouin Zone (GBZ). The non-Bloch entanglement entropy is defined on a quasi-reciprocal lattice, obtained by performing an ordinary Fourier transformation on the non-Bloch Hamiltonian. We demonstrate that the broken bulk-boundary correspondence is recovered in terms of the non-Bloch entanglement entropy. When the GBZ is circular, we show that the non-Bloch entanglement entropy is well-defined (real and positive-definite) in large parameter regions, except close to the exceptional points (EPs). In the critical region, we found that each Fermi point contributes precisely 1 to the central charge $c$ of the logarithmic scaling. At the EP, the central charge becomes negative due to the presence of the exceptional bound state. For the case of non-circular GBZ, long-range hopping emerges in the quasi-reciprocal lattice, and the von Neumann entropy on the GBZ is no longer real. However, the non-Bloch edge entanglement entropy remains real, which serves as a reliable topological indicator and respects the bulk-boundary correspondence. We compute the topological phase diagram and reveal the critical behavior along the exceptional phase boundaries.
@@
Non-Hermitian quantum mechanics naturally arises from the study of open quantum systems coupling with an external environment [1, 2]. For non-Hermitian systems with nonreciprocal hoppings, the bulk eigenstates under open boundary conditions (OBC) accumulate at boundaries, which is named by the non-Hermitian skin effect (NHSE) [3, 4, 5, 6] and has been extensively studied recently [7, 8, 9, 10]. In the presence of NHSE, the energy spectrum sensitively depends on the boundary conditions: the energy spectrum under OBC forms a star-shaped branching figures [11], while the energy spectrum under periodic boundary conditions (PBC) makes up loops enclosing the OBC spectrum [12, 13, 14]. The NHSE also results in the breakdown of the celebrated bulk-boundary correspondence (BBC) of topological band theory [15, 16]: the transition points of bulk topological invariants do not match with transition points of zero modes under OBC [17, 18, 3]. In order to understand the topology of non-Hermitian systems, the non-Bloch band theory and the generalized Brillouin zone (GBZ) are proposed, based on which the BBC for non-Hermitian systems can be reestablished [3, 19, 20]. Motivated by this breakthrough, the non-Bloch band theory has been widely used, e.g., to construct Green function [21], to investigate the dynamical phenomena [22] and the non-Bloch parity-time symmetry [23].
@@
The entanglement properties of the non-Hermitian quantum systems is also an interesting topic which has attracted much attentions recently [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. The conception of EE has been generalized to the non-Hermitian systems in the biorthogonal basis [32, 31]. A logarithmic scaling with central charge $c=-2$ is found at critical point of the non-Hermitian Su-Schrieffer-Heeger (SSH) model, which can be explained by the $bc$-ghost non-unitary CFT [32, 41]. However, the physical meaning of EE remains somewhat ambiguous in this straightforward generalization [41, 43]: The eigenvalues of reduced density matrix $\rho_{A}$ are no longer positive definite, which leads to a negative or even complex EE. This is inconsistent with the probability interpretation of a measurable quantity [41, 43]. Some attempts have been made to remedy this situation, for example, the generic entanglement entropy $S_{A}=-\mathrm{Tr}(\rho_{A}\ln|\rho_{A}|)$ is proposed in Ref. [41], the modified partial trace formalism is introduced in Ref. [30]. However, more systematic analysis is needed in this direction.
Inspired by the success of non-Hermitian topological band theory in terms of GBZ, we may ask whether it is possible to construct an EE on the GBZ that has proper behaviors? We address this question in this paper for the non-Hermitian systems with skin effect. We begin with a GBZ Hamiltonian and then make a Fourier transformation to obtain a quasi-reciprocal lattice [58]. The entanglement entropy on GBZ is calculated by partitioning this quasi-reciprocal lattice. The behaviors of various EE will be examined for two types of GBZ (circular or non-circular). Our study may shed light on the physical meanings of GBZ and non-Hermitian EE, and also open the possibility to study the many-body physics based on GBZ.
This paper is organized as follows: In Sec. II, we review the basic notions of biorthogonal entanglement entropy. In Sec. III, we introduce the non-Hermitian SSH model and the quasi-reciprocal lattice, on which the entanglement entropy is calculated. In Sec. IV, we present the results of entanglement entropy on the circular GBZ. In Sec. V, we present the results of entanglement entropy on the non-circular GBZ. Finally, a brief summary and discussion are provided in Sec. VI.
@@
in which the sites $i,j$ are restricted inside the subsystem $A$, and the index $m$ labels the occupied energy levels. The single-particle entanglement spectrum is obtained from the eigenvalue spectrum $\{\xi_{l}\}$ of the matrix $C^{A}$ under periodic boundary condition. Based on this, the von-Neumann and Rényi entropy are given by
@@
where $S^{\alpha}_{\text{OBC}}$ and $S^{\alpha}_{\text{PBC}}$ are $\alpha$-th Rényi entropies calculated under OBC and PBC, respectively. Here, half of $S_{PBC}^{\alpha}$ is subtracted to eliminate the leading terms of entanglement. Note that, in defining entanglement, there are two cuts in the case of PBC but only one cut in the case of OBC [64, 33].
The non-Hermitian entanglement entropy is a useful tool for detecting topological phase transitions and extracting information about edge states [64, 33, 60]. In addition, it can reveal critical behavior within a system [32, 39, 40]: by examining the scaling of entanglement entropy with subsystem size, valuable information about critical points and exceptional points can be obtained. The entanglement entropy can be fitted using the following universal formula [65, 66],
@@ III Quasi-reciprocal Hamiltonian based on generalized Brillouin zone
in which $k\in[0,2\pi)$ is the crystal momentum. The hopping processes of this model in real space are illustrated in Fig. 1(a). According to the theory of GBZ [3, 19, 20], we substitute $e^{ik}$ with $\beta=e^{ik^{\prime}}$ ($k^{\prime}=k+i\tau(k)\in\mathbb{C}$) to gain the corresponding non-Bloch Hamiltonian
@@
The condition (15) implies that the two exponential eigenstates, corresponding to $\beta_{2}$ and $\beta_{3}$, are of the same order. This allows them to cancel each other near the boundaries of an open lattice, forming a standing wave that satisfies the OBC [3, 19]. Two typical GBZs are shown in Fig. 2(a) and (b), where the GBZ forms a closed loop encircling the origin on the complex plane. The OBC spectrum in the thermodynamic limit can be exactly obtained when $\beta$ goes along the loop of GBZ.
To define the entanglement entropy in terms of GBZ, one needs to construct an artificial real-space lattice which respects the translation invariance. Suppose that the phase angle of $\beta_{GBZ}$ is equally distributed on the GBZ, we then construct non-Bloch Hamiltonian matrix $H(\beta_{GBZ})$ by selecting the points whose phase angle belongs to the Brillouin zone (BZ) of original physical lattices, that is,
@@
where $N$ is the number of unit cells of the lattice ($N=L/2$). The non-Bloch operator $c_{\beta(k_{m})}^{\dagger}$ can be transformed into a new lattice operator $\tilde{c}_{j}$ by a Fourier transformation,
@@
Figure 2: The examples of circular GBZ (a) and non-circular GBZ (b). The green dashed circle denotes the conventional BZ with $|\beta|=1$. The radius of circular GBZ (orange circle) in (a) is everywhere smaller than $1$, indicating that wave functions only accumulate at the left end of the chain. Sub-figures (c) and (d) show the energy spectra of original OBC Hamiltonian $H_{OBC}$ (blue dots) and quasi-reciprocal OBC Hamiltonian $\widetilde{H}_{OBC}$ (red circles). The length of lattice is set to be $L=40$. For the case of circular GBZ (c), the blue dots precisely match the red circles. However, for the non-circular GBZ (d), the blue dots slightly deviate from the red circles, especially close to the bifurcations. The inset of sub-figure (d) shows the mean square error (vertical axis) between energy spectra of $H_{OBC}$ and $\widetilde{H}_{OBC}$ as a function of system length $L$ (horizontal axis). We fix $t_{3}=0$ in (a) and (c), but $t_{3}=0.2$ in (b) and (d). Other parameters used in the calculations are $t_{1}=0.4$, $t_{2}=1$, and $\gamma=2/3$.
@@
$\{\epsilon_{OBC}\}\stackrel{{\scriptstyle L\to\infty}}{{=}}\{\widetilde{\epsilon}_{OBC}\}\stackrel{{\scriptstyle L\to\infty}}{{=}}\{\widetilde{\epsilon}_{PBC}\}$ (19)
When next-nearest-neighbor (NNN) hoppings are included in the non-Hermitian SSH model, the degree of polynomial $\det[E-H(\beta)]$ is greater than 2, and the GBZ will become non-circular [3]. In this case, the module $|\beta|$ is not a constant but depends on the phase angle $k$. By Fourier transformation, this leads to the long-range hoppings in quasi-reciprocal lattices [58]. In practical calculations, we adopt an approximation to obtain the open boundary of quasi-reciprocal lattice: we neglect all the hopping terms whose hopping range is larger than half of the lattice length ($L/2$). This is reasonable since long-range hoppings are generically power-law decaying [58].
@@
with $\tilde{S}=\mathrm{diag}\{r,r,r^{2},r^{2},\cdots,r^{N-1},r^{N},r^{N}\}$. This transformation differs from the well-studied similarity transformation $\bar{H}_{OBC}=\bar{S}^{-1}H_{OBC}\bar{S}$ in Ref. [3] with $\bar{S}=\mathrm{diag}\{1,r,r,r^{2},\cdots,r^{N-1},r^{N}\}$, which renormalizes the intracell hopping amplitudes; see Fig. 1(c). Note that $\widetilde{H}_{OBC}$ is also connected to $\bar{H}_{OBC}$ by a similarity transformation, $\widetilde{H}_{OBC}=S^{-1}\bar{H}_{OBC}S$, with $S=\mathrm{diag}\{r,1,r,1,\cdots,r,1\}$. Therefore, $\widetilde{H}_{OBC}$ and $\bar{H}_{OBC}$ have the same energy spectra and entanglement entropy.
If $t_{3}\neq 0$, the long-range hoppings appear in the quasi-reciprocal lattices due to the non-constant $|\beta|$. There are tiny difference between the energy spectra of $H_{OBC}$ and $\widetilde{H}_{OBC}$ [see Fig. 2(d)], for the long-range hoppings can not be fully taken into account when the length of lattices is finite. The difference in energy spectra will decrease to zero in the thermodynamic limit.
@@ IV Entanglement entropy on Circular GBZ
We first present the results of circular GBZ in Fig. 3, in which we use the non-Hermitian SSH model with $t_{3}=0$ as an example. It is well-known that the BBC is broken in the non-Hermitian SSH model: the topological transition of zero-energy modes under OBC can not be explained by the Bloch Hamiltonian $h(k)$ under PBC [3, 19]; see Fig. 3(a). The entanglement entropy can be calculated by applying the correlation matrix method described in Sec. II to the original non-Hermitian Hamiltonian under PBC [31, 32]. We show the results of entanglement entropy for SSH model with $t_{3}=0$ in Fig. 3(b). The entanglement entropy is well-defined in the line-gap region, i.e., the values are real and positive-definite (away from the two exceptional points). However, in the point-gap region, the entanglement entropy is not well-defined: it becomes complex with a non-zero imaginary part [31, 60]. Note that at the special point with $t_{1}=t_{2}$ in the point-gap region, the entanglement entropy is also real, as studied in Ref. [39]. In all, the entanglement entropy defined on the original PBC lattice basically reflects the characteristics of the Bloch band, and does not respect the BBC.
@@
The entanglement entropy drops to $2\log(2)$ at the point $t_{1}=\gamma$; see Fig. 3(c). At this point, the GBZ collapses into a single point with $r=0$, and the intra-(inter-)unit cell hoppings of the quasi-reciprocal lattice become $2\gamma$ and 0 (0 and $\infty$), respectively. By performing the similar transformation $S=\mathrm{diag}\{1,1/r,1,1/r,\cdots,1,1/r\}$, this case can be mapped into the dimerized limit with intra-unit cell hopping being zero; see Fig. 1(c). There are two isolated edge states, hence the entanglement entropy is exactly $2\log(2)$ [68]. Note, when $t_{1}<\gamma$, the Hamiltonian $H$ cannot be transformed into a Hermitian SSH model [3]. However, the entropy remains well-defined, as demonstrated in Fig. 3(c); see the Appendix A for the analysis of realness. This is a significant finding of the present paper.
@@
We now investigate the critical behavior of the quasi-reciprocal lattice Hamiltonian, which exhibits an EP at $\gamma_{c}=\sqrt{t_{1}^{2}+t_{2}^{2}}$. We firstly examine the region away from the EP ($\gamma\neq\gamma_{c}$). In Fig. 4(a), the fitted central charge $c$ is plotted as a function of $\gamma$ for $t_{1}=0.5$. For $\gamma<\gamma_{c}$, the value of $c$ is zero, indicating that the system is noncritical in this region. However, for $\gamma>\gamma_{c}$, the value of $c$ saturates to 2, implying critical behavior; See Fig. 4(b) for the representative logarithmic scaling of entanglement entropy with $L_{A}/L=1/2$. The critical behavior is associated with the Fermi surface structure of the energy spectrum [39]. Since we construct the ground state according to the real part of energy, the system is an insulator without any Fermi points when $\gamma<\gamma_{c}$, but when $\gamma>\gamma_{c}$ it becomes a gapless metal with two Fermi points. We therefore conclude that each Fermi point contributes exactly 1 to the central charge $c$ of the logarithmic scaling. This is different from the original SSH model (PBC) studied in Ref. [39], where each Fermi point contributes 1/2 to the value of $c$.
@@
At the EP with $\gamma=\gamma_{c}$, the entanglement entropy exhibits divergent behavior. Please refer to the inset in Fig. 4(c) for the exceptional momentum $k_{EP}$ (0 and $\pi$) in this model, and note that the dispersion of these EPs is square root dependence on $k$ instead of linear. To avoid divergence, the entanglement entropy can be calculated by slightly shifting from the EP $\gamma=\gamma_{c}-\eta$ [32, 41]. The scaling behavior of entanglement entropy for a fixed system length $L=400$ is shown in Fig. 4(d), where a tiny shift $\eta=1\times 10^{-8}$ is used and $L_{A}$ is chosen to include an odd number of unit cells. The extracted central charge is negative (around $-3$), which is attributed to the exceptional bound state close to EP [40]. However, the value of central charge slightly increases when the system length is increasing, as depicted in Fig. 4(c).
@@
Figure 4: (a) The central charge $c$ of entanglement entropy on circular GBZ as a function of $\gamma$, excluding the EP at $\gamma=\gamma_{c}$. The two insets display representative energy spectra on two sides of the EP $\gamma_{c}=\sqrt{1.25}$, where the solid (dashed) line indicate the occupied (unoccupied) states. (b) The logarithmic scaling of entanglement entropy (fixed ratio $L_{A}/L=1/2$) when $\gamma$ is far away from EP. (c) The fitted central charge at the EP as a function of total system length $L$. In the calculations, $\gamma$ is shifted from $\gamma_{c}$ by a tiny quantity $\eta=1\times 10^{-8}$, and the formula in Eq. (10) is used for the fitting process. Insets display the energy spectrum and the dispersion relation, revealing two exceptional momenta at $k=0,\pi$ that render the momentum-space Hamiltonian defective. (d) The entanglement entropy (green circles) as a function of the subsystem length $L_{A}$, for a fixed length $L=400$ and at $\gamma=\gamma_{c}-\eta$. The black line is the fitting curve of entanglement entropy according to Eq. (10), from which the central charge can be extracted. The hopping parameters for all sub-figures are $t_{1}=0.5$, $t_{2}=1$, and $t_{3}=0$.